A train carries 92 passengers altogether. Some cars carry 6 passengers and some cars carry 4 passengers. There are three less 6 passenger cars than 4 passengers cars. How many 4 passenger cars are there?
Answer by stanbon(75887)

A train carries 92 passengers altogether. Some cars carry 6 passengers and some cars carry 4 passengers. There are three less 6 passenger cars than 4 passengers cars. How many 4 passenger cars are there?
-------------
Equations:
6s + 4f = 92 passengers
-----
s = f-3
-----------------
Substitute for "s" and solve for "f":
6(f-3) + 4f = 92
6f-18 + 4f = 92
10f = 110
f = 11 (# of four-passenger cars)
===================================
Cheers,
Stan H.
===========